All the Lucida text fonts

prepared by Sebastian Rahtz

February 19th 1995

hlx/b/it hlxdi8t (LucidaFax-DemiItalic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlx/b/n hlxd8t (LucidaFax-Demi): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlx/m/it hlxrir8t (LucidaFax-Italic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlx/m/n hlxr8t (LucidaFax): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlh/b/it hlcdib8t (LucidaBright-DemiItalic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlh/b/n hlcdb8t (LucidaBright-Demi): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlh/m/it hlcrib8t (LucidaBright-Italic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlh/m/n hlcrb8t (LucidaBright): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlce/m/it hlcrie8t (LucidaCalligraphy-Italic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés hlcf/m/n hlcrf8t (LucidaBlackletter): Animadversion for a giraffe costs £123.

Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlcn/m/it hlcrin8t (LucidaCasual-Italic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlcn/m/n hlcrn8t (LucidaCasual): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlst/b/n hlsbt8t (LucidaSans-TypewriterBold): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlst/b/sl hlsbot8t (LucidaSans-TypewriterBoldOblique): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hls/ub/it hlsbi8t (LucidaSans-BoldItalic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hls/ub/n hlsb8t (LucidaSans-Bold): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hls/b/it hlsdi8t (LucidaSans-DemiItalic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hls/b/n hlsd8t (LucidaSans-Demi): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hls/m/it hlsri8t (LucidaSans-Italic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hls/m/n hlsr8t (LucidaSans): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlct/b/n hlcbt8t (LucidaTypewriterBold): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés hlct/b/sl hlcbot8t (LucidaTypewriterOblique): Animadversion for

a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés

hlcw/m/it hlcriw8t (LucidaHandwriting-Italic): Animadversion for a giraffe costs £123. Waß ist das für ein Klopf? We are often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés